OpenStudy (anonymous):

Use inductive reasoning to find the next term in the sequence 0, 1, 3, 6, 10… What is the next term? A. 14 B. 15 C. 16 D. 26

OpenStudy (anonymous):

@JCMMusic

OpenStudy (anonymous):

I got it. :) 0+1=1 1+2=3 3+3=6 6+4=10 10+5=15

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

The next term is 15.

OpenStudy (anonymous):

@OrangeMaster Awesome! :)